# Generalized Neural Sorting Networks with Error-Free Differentiable Swap Functions

### v1

@@ Appendix D Proof of Proposition 1
If two elements at indices $i$ and $j$ are swapped by a single swap function, $[\mathbf{P}]_{kk}=1$ for $k\in[n]\backslash\{i,j\}$, $[\mathbf{P}]_{kl}=0$ for $k\neq l$, $k,l\in[n]\backslash\{i,j\}$, $[\mathbf{P}]_{ii}=[\mathbf{P}]_{jj}=p$, and $[\mathbf{P}]_{ij}=[\mathbf{P}]_{ji}=1-p$, where $p$ is the output of sigmoid function, i.e., $p=\sigma(y-x)$ or $p=\lfloor\sigma(y-x)\rceil$. Since the multiplication of doubly-stochastic matrices is still doubly-stochastic, Proposition 1 is true. ∎
@@ Appendix E Proof of Proposition 2
Let $\overline{\mathrm{min}}^{k}(x,y)$ and $\overline{\mathrm{min}}^{k}(x,y)$ be minimum and maximum values where $\mathrm{swap}$ with $\overline{\mathrm{min}}$ and $\overline{\mathrm{max}}$ is applied $k$ times repeatedly. By Definition 1, the following inequality is satisfied:
$\min(x,y)<\overline{\mathrm{min}}^{1}(x,y)<\cdots<\overline{\mathrm{min}}^{k}(x,y)\leq\overline{\mathrm{max}}^{k}(x,y)<\cdots<\overline{\mathrm{max}}^{1}(x,y)<\max(x,y),$ (21)
under the assumption that $\nabla_{x}\sigma(x)>0$. If $k\to\infty$, $\overline{\mathrm{min}}^{k}(x,y)=\overline{\mathrm{max}}^{k}(x,y)$. Therefore, a softening error is $(\max(x,y)-\min(x,y))/2$ since $\overline{\mathrm{max}}^{k}(x,y)=(\min(x,y)+\max(x,y))/2$ by (9). Note that the assumption $\nabla_{x}\sigma(x)>0$ implies that $\sigma(\cdot)$ is a strictly monotonic sigmoid function. ∎
@@ Appendix F Proof of Proposition 3
According to Definition 1, given $x$ and $y$, the softening error $x^{\prime}-\min(x,y)$ is expressed as
@@
Appendix G Split Strategy to Reduce the Number of Possible Permutations
Figure 7: A split process of $\mathbf{P}$
As presented in (14) and Proposition 1, the permutation matrix for the error-free DSF is a discretized doubly-stochastic matrix, which can be denoted as $\mathbf{P}_{\mathrm{hard}}$, in a forward pass, and is differentiable in a backward pass. Here, we show an interesting proposition of $\mathbf{P}_{\mathrm{hard}}$:
@@ Appendix G Split Strategy to Reduce the Number of Possible Permutations
Let $\mathbf{s}\in\mathbb{R}^{n}$ and $\mathbf{P}_{\mathrm{hard}}\in\mathbb{R}^{n\times n}$ be an unordered sequence and its corresponding permutation matrix to transform to $\mathbf{s}_{\mathrm{o}}$, respectively. We are able to split $\mathbf{s}$ to two subsequences $\mathbf{s}_{1}\in\mathbb{R}^{n_{1}}$ and $\mathbf{s}_{2}\in\mathbb{R}^{n_{2}}$ where $\mathbf{s}_{1}=[\mathbf{s}]_{1:n_{1}}$ and $\mathbf{s}_{2}=[\mathbf{s}]_{n_{1}+1:n_{1}+n_{2}}$. Then, $\mathbf{P}_{\mathrm{hard}}$ is also split to $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ and $\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$, so that $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are (discretized) doubly-stochastic.
@@
Since a split does not change the relative order of elements in the same split and entries in the permutation matrix is zero or one, a permutation matrix can be split as shown in Figure 7. Moreover, multiple splits are straightforwardly available. ∎
In contrast to $\mathbf{P}_{\mathrm{hard}}$, it is impossible to split $\mathbf{P}_{\mathrm{soft}}$ to sub-block matrices since such sub-block matrices cannot satisfy the property of doubly-stochastic matrix, which is discussed in Proposition 1. Importantly, Proposition 4 does not show a possibility of the recoverable decomposition of the permutation matrix, which implies that we cannot guarantee the recovery of decomposed matrices to the original matrix. Regardless of the existence of recoverable decomposition, we intend to reduce the number of possible permutations with sub-block matrices, rather than holding the large number of possible permutations with the original permutation matrix. Therefore, by Proposition 4, relative relationships between instances with a smaller number of possible permutations are more distinctively learnable than the relationships with a larger number of possible permutations, preventing a sparse correct permutation among a large number of possible permutations.
Appendix H Details of Architectures
We describe the details of the neural architectures used in our paper, as shown in Tables 5, 6, 7, 8, 9, 10, 11, 12, 13, and 14. For the experiments of image fragments, we omit some of the architectures with particular fragmentation, since they follow the same architectures presented in Tables 11, 12, 13, and 14. Only differences are the size of inputs, and therefore the respective sizes of the first fully-connected layers are changed.
Table 5: Architecture of the convolutional neural networks for the four-digit MNIST dataset
@@ Appendix H Details of Architectures
Table 6: Architecture of the Transformer-Small models for the four-digit MNIST dataset
@@
Table 7: Architecture of the Transformer-Large models for the four-digit MNIST dataset
@@
Table 8: Architecture of the convolutional neural networks for the SVHN dataset
@@
Table 9: Architecture of the Transformer-Small models for the SVHN dataset
@@
Table 10: Architecture of the Transformer-Large models for the SVHN dataset
@@
Table 11: Architecture of the convolutional neural networks for the MNIST dataset of 4 image fragments of size $14\times 14$
@@
Table 12: Architecture of the Transformer models for the MNIST dataset of 4 image fragments of size $14\times 14$
@@
Table 13: Architecture of the convolutional neural networks for the CIFAR-10 dataset of 4 image fragments of size $16\times 16$
@@
Table 14: Architecture of the Transformer models for the CIFAR-10 dataset of 4 image fragments of size $16\times 16$
@@
Appendix I Details of Experiments
As described in the main article, we use three public datasets: MNIST (LeCun et al., 1998), SVHN (Netzer et al., 2011), and CIFAR-10 (Krizhevsky & Hinton, 2009). Unless otherwise specified, the learning rate $10^{-3.5}$ is used for the CNN architectures and the learning rate $10^{-4}$ is used for the Transformer-based architectures; see the scripts for the exact learning rates we use in the experiments. Learning rate decay is applied to multiply $0.5$ in every 50,000 steps for the experiments on sorting multi-digit images and every 20,000 steps for the experiments on sorting image fragments. Moreover, we balance two objectives for $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$ by multiplying $1$, $0.1$, $0.01$, or $0.001$; see the scripts included in our implementation for the respective values for all the experiments. For random seeds, we pick five random seeds for all the experiments: $42$, $84$, $126$, $168$, and $210$; these values are picked without any trials. Other missing details can be found in our implementation, which is included alongside our manuscript. Furthermore, we use several commercial NVIDIA GPUs, i.e., GeForce GTX Titan Xp, GeForce RTX 2080, and GeForce RTX 3090, for the experiments.
Appendix J Study on Balancing Hyperparameter
We conduct a study on a balancing hyperparameter in the experiments on sorting the four-digit MNIST dataset, as shown in Table 15. For these experiments, we use steepness 6, 20, 29, 32, 25, and 124 for sequence lengths 3, 5, 7, 9, 15, and 32, respectively. Also, we use a learning rate $10^{-3}$ and five random seeds 42, 84, 126, 168, and 210.
Table 15: Study on a balancing hyperparameter $\lambda$ in the experiments of sorting the four-digit MNIST dataset
@@ Appendix I Details of Experiments
1.000 94.9 (96.5) 86.1 (93.8) 73.7 (90.6) 60.8 (87.9) 10.8 (67.8) 0.0 (31.8)
0.100 94.8 (96.4) 86.4 (93.8) 75.3 (91.2) 65.6 (89.6) 33.7 (83.0) 1.2 (63.0)
0.010 94.9 (96.5) 87.2 (94.2) 76.3 (91.6) 65.6 (89.5) 32.9 (82.8) 0.8 (62.6)
0.001 95.2 (96.7) 86.6 (94.0) 75.3 (91.2) 64.6 (89.1) 34.4 (83.2) 1.4 (66.4)
0.000 94.8 (96.4) 86.6 (93.9) 76.4 (91.7) 64.0 (89.0) 33.0 (82.6) 1.3 (66.0)
Appendix K Study on Steepness and Learning Rate
We present studies on steepness and learning rate for the experiments on sorting the multi-digit MNIST dataset, as shown in Tables 16, 17, 18, 19, 20, and 21. For these experiments, a random seed 42 is only used due to numerous experimental settings. Also, we use balancing hyperparameters $\lambda$ as 1.0, 1.0, 0.1, 0.1, 0.1, and 0.1 for sequence lengths 3, 5, 7, 9, 15, and 32, respectively. Since there are many configurations of steepness, learning rate, and a balancing hyperparameter, we cannot include all the configurations here. The final configurations we use in the experiments are described in our implementation.
@@
Appendix L Limitations
Because our methods are based on a sorting network, our ideas are limited to sorting algorithms. It implies that it is not easy to devise a neural network-based approach to solving general problems in computer science, e.g., combinatorial optimization, with our ideas.
Appendix M Broader Impacts
Since our work proposes a method to sort high-dimensional instances in a sequence using a differentiable swap function, a direct negative societal impact does not exist. However, there can be a potential issue of sorting controversial high-dimensional data that is considered not to be sorted, e.g., beauty and intelligence.

### v2

@@ Appendix D Proof of Proposition 1
If two elements at indices $i$ and $j$ are swapped by a single swap function, $[\mathbf{P}]_{kk}=1$ for $k\in[n]\backslash\{i,j\}$, $[\mathbf{P}]_{kl}=0$ for $k\neq l$, $k,l\in[n]\backslash\{i,j\}$, $[\mathbf{P}]_{ii}=[\mathbf{P}]_{jj}=p$, and $[\mathbf{P}]_{ij}=[\mathbf{P}]_{ji}=1-p$, where $p$ is the output of a sigmoid function, i.e., $p=\sigma(y-x)$ or $p=\lfloor\sigma(y-x)\rceil$. Since the multiplication of doubly-stochastic matrices is still doubly-stochastic, Proposition 1 is true. ∎
@@ Appendix E Proof of Proposition 2
Let $\overline{\mathrm{min}}^{k}(x,y)$ and $\overline{\mathrm{min}}^{k}(x,y)$ be minimum and maximum values where $\mathrm{swap}$ with $\overline{\mathrm{min}}$ and $\overline{\mathrm{max}}$ is applied $k$ times repeatedly. By Definition 1 and $\overline{\mathrm{min}}^{i}<\overline{\mathrm{min}}^{i+1}$ and $\overline{\mathrm{max}}^{i+1}<\overline{\mathrm{max}}^{i}$, the following inequalities are satisfied:
$\min(x,y)<\overline{\mathrm{min}}^{1}(x,y)<\overline{\mathrm{min}}^{2}(x,y)<\cdots<\overline{\mathrm{min}}^{k}(x,y)$
$\leq\overline{\mathrm{max}}^{k}(x,y)<\cdots<\overline{\mathrm{max}}^{2}(x,y)<\overline{\mathrm{max}}^{1}(x,y)<\max(x,y),$ (21)
under the assumption that $\nabla_{x}\sigma(x)>0$. By (21) and $\nabla_{x}\sigma(x)>0$, we can obtain the following inequality:
$0\leq\overline{\mathrm{max}}^{k+1}(x,y)-\overline{\mathrm{min}}^{k+1}(x,y)<\overline{\mathrm{max}}^{k}(x,y)-\overline{\mathrm{min}}^{k}(x,y)<\overline{\mathrm{max}}^{k-1}(x,y)-\overline{\mathrm{min}}^{k-1}(x,y).$ (22)
Therefore, $\lim_{k\to\infty}\overline{\mathrm{max}}^{k}(x,y)-\overline{\mathrm{min}}^{k}(x,y)=0$, and $\overline{\mathrm{min}}^{k}(x,y)=\overline{\mathrm{max}}^{k}(x,y)$ if $k\to\infty$. To sum up, a softening error for $k\to\infty$ is $(\max(x,y)-\min(x,y))/2$ since $\overline{\mathrm{max}}^{k}(x,y)=(\min(x,y)+\max(x,y))/2$ by (9). Note that the assumption $\nabla_{x}\sigma(x)>0$ implies that $\sigma(\cdot)$ is a strictly monotonic sigmoid function. ∎
@@ Appendix F Proof of Proposition 3
According to Definition 1, given $x$ and $y$, the softening error $x^{\prime}-\min(x,y)$ is expressed as the following:
@@
Appendix G Details of Permutation-Equivariant Networks with Multi-Head Attention
Figure 7: Simple illustration of our permutation-equivariant network with multi-head attention. For the sake of brevity, we do not depict the details of multi-head attention, layer normalization, and feed-forward networks. Note that FE stands for a feature extractor.
Figure 7 illustrates the Transformer-based permutation-equivariant network, which is implemented with multi-head attention (Vaswani et al., 2017). For the sake of brevity, this illustration briefly depicts our permutation-equivariant network without detailing the specifics of the Transformer network. Each instance in a sequence is first processed by a feature extractor, i.e., a convolutional neural network. Then, a sequence of latent vectors is provided into the Transformer network without positional encoding. At each multi-head attention module, each latent vector is updated by the aggregation of the latent vectors given where the aggregation is determined by the operations explained in Section 4 and (16). After passing through multiple layers of multi-head attention and the corresponding components such as layer normalization and feed-forward neural networks, the final fully-connected layer is applied to transform the outputs of the Transformer network into a score vector $\mathbf{s}$. The details of the Transformer network can be found in the work by Vaswani et al. (2017).
Appendix H Split Strategy to Reduce the Number of Possible Permutations
Figure 8: A split process of $\mathbf{P}$.
As presented in (14) and Proposition 1, the permutation matrix for the error-free DSF is a discretized doubly-stochastic matrix, which is denoted as $\mathbf{P}_{\mathrm{hard}}$, in a forward pass, and is differentiable in a backward pass. Here, we show an interesting proposition of $\mathbf{P}_{\mathrm{hard}}$:
@@ Appendix H Split Strategy to Reduce the Number of Possible Permutations
Let $\mathbf{s}\in\mathbb{R}^{n}$ and $\mathbf{P}_{\mathrm{hard}}\in\mathbb{R}^{n\times n}$ be an unordered sequence and the corresponding permutation matrix to transform it to $\mathbf{s}_{\mathrm{o}}$, respectively. We are able to split $\mathbf{s}$ to two subsequences $\mathbf{s}_{1}\in\mathbb{R}^{n_{1}}$ and $\mathbf{s}_{2}\in\mathbb{R}^{n_{2}}$ where $\mathbf{s}_{1}=[\mathbf{s}]_{1:n_{1}}$ and $\mathbf{s}_{2}=[\mathbf{s}]_{n_{1}+1:n_{1}+n_{2}}$. Then, $\mathbf{P}_{\mathrm{hard}}$ is also split to $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ and $\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$, so that $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are (discretized) doubly-stochastic.
@@
A split does not change the relative order of elements in the same split and each entry in the permutation matrix is zero or one, so that a permutation matrix can be split as shown in Figure 8. Moreover, multiple splits are straightforwardly doable. ∎
In contrast to $\mathbf{P}_{\mathrm{hard}}$, it is impossible to split $\mathbf{P}_{\mathrm{soft}}$ to sub-block matrices since such sub-block matrices cannot satisfy the property of doubly-stochastic matrix, which is discussed in Proposition 1. Importantly, Proposition 4 does not show a possibility of the recoverable decomposition of the permutation matrix, which implies that we cannot guarantee the recovery of decomposed matrices to the original matrix. Regardless of the existence of recoverable decomposition, we attempt to reduce the number of possible permutations with sub-block matrices, rather than holding the large number of possible permutations with the original permutation matrix. Therefore, by Proposition 4, relative relationships between instances with a smaller number of possible permutations are more distinctively learnable than the relationships with a larger number of possible permutations, preventing a sparse correct permutation among a large number of possible permutations.
Appendix I Details of Architectures
We describe the details of the neural architectures used in our paper, as shown in Tables 5, 6, 7, 8, 9, 10, 11, 12, 13, and 14. For the experiments on sorting image fragments, we omit some of the architectures employed for particular fragmentation, since they follow the same architectures presented in Tables 11, 12, 13, and 14. Only differences are the sizes of inputs, and therefore the respective sizes of the first fully-connected layers change.
Table 5: Architecture of the convolutional neural networks for the four-digit MNIST dataset.
@@ Appendix I Details of Architectures
Table 6: Architecture of the Transformer-Small models for the four-digit MNIST dataset.
@@
Table 7: Architecture of the Transformer-Large models for the four-digit MNIST dataset.
@@
Table 8: Architecture of the convolutional neural networks for the SVHN dataset.
@@
Table 9: Architecture of the Transformer-Small models for the SVHN dataset.
@@
Table 10: Architecture of the Transformer-Large models for the SVHN dataset.
@@
Table 11: Architecture of the convolutional neural networks for the MNIST dataset of 4 image fragments of size $14\times 14$.
@@
Table 12: Architecture of the Transformer models for the MNIST dataset of 4 image fragments of size $14\times 14$.
@@
Table 13: Architecture of the convolutional neural networks for the CIFAR-10 dataset of 4 image fragments of size $16\times 16$.
@@
Table 14: Architecture of the Transformer models for the CIFAR-10 dataset of 4 image fragments of size $16\times 16$.
@@
Appendix J Details of Experiments
As described in the main article, we use three public datasets: MNIST (LeCun et al., 1998), SVHN (Netzer et al., 2011), and CIFAR-10 (Krizhevsky & Hinton, 2009). Unless otherwise specified, a learning rate $10^{-3.5}$ is used for the CNN architectures and a learning rate $10^{-4}$ is used for the Transformer-based architectures; see our implementation for the exact learning rates we utilize in the experiments. Learning rate decay is applied by multiplying $0.5$ in every 50,000 steps for the experiments on sorting multi-digit images and every 20,000 steps for the experiments on sorting image fragments. Moreover, we balance two objectives for $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$ by multiplying $1$, $0.1$, $0.01$, or $0.001$; see our implementation for the respective values for all the experiments. For random seeds, we pick five random seeds $42$, $84$, $126$, $168$, and $210$ for all the experiments; these values are picked without any trials. Other missing details can be found in our implementation. Furthermore, we employ several commercial NVIDIA GPUs, i.e., GeForce GTX Titan Xp, GeForce RTX 2080, and GeForce RTX 3090, in the experiments.
Appendix K Study on Balancing Hyperparameter
Table 15: Study on a balancing hyperparameter $\lambda$ in the experiments on sorting the four-digit MNIST dataset.
@@
1.000 94.8 (96.4) 86.9 (94.1) 74.2 (90.9) 62.6 (88.6) 12.0 (69.4) 0.0 (38.9)
0.100 94.9 (96.5) 87.2 (94.2) 75.3 (91.3) 64.7 (89.2) 34.7 (83.3) 2.1 (69.2)
0.010 95.1 (96.6) 87.1 (92.7) 75.2 (91.2) 63.5 (88.8) 33.2 (82.7) 0.7 (60.7)
0.001 95.2 (96.7) 87.0 (94.1) 76.6 (91.6) 64.8 (89.2) 32.9 (82.7) 0.5 (61.5)
0.000 94.9 (96.5) 87.2 (94.2) 75.9 (91.5) 64.4 (89.1) 34.1 (83.2) 0.9 (60.6)
We conduct a study on a balancing hyperparameter in the experiments on sorting the four-digit MNIST dataset, as shown in Table 15. For these experiments, we use steepness 2, 14, 23, 38, 25, and 124 for sequence lengths 3, 5, 7, 9, 15, and 32, respectively. Also, we use a learning rate $10^{-3}$ and 5 random seeds 42, 84, 126, 168, and 210.
Appendix L Study on Steepness and Learning Rate
[…]
@@
Appendix M Limitations
While a sorting task is one of the most significant problems in computer science and mathematics (Cormen et al., 2022), our ideas, which are built on sorting networks (Knuth, 1998; Ajtai et al., 1983), can be limited to sorting algorithms. It implies that it is not easy to devise neural network-based approaches to solving general problems in computer science, e.g., combinatorial optimization, which are inspired by our ideas.
In addition, while our proposed methods show the superior performance compared to the baseline methods, this line of research suffers from performance degradation for longer sequences as shown in Tables 12, and 3. More precisely, for longer sequences, the element-wise accuracy does not decline dramatically, but the sequence-wise accuracy significantly drops due to the nature of sequences. Incorporating our contributions such as the error-free DSFs and the Transformer-based networks, we expect that the further progress of neural network-based sorting networks can be achieved. In particular, the consideration of more sophisticated neural networks, which are capable of handling longer sequences, might help improve performance. This will be left for future work.
Our frameworks successfully learn relationships between high-dimensional data with ordinal contents as shown in Section 5. However, we suppose that our methods might fail in sorting data without ordinal information; the elaborate discussion on this topic can be found in Section 7. In order to sort more ambiguous high-dimensional data, we can combine our work with part-based or segmentation-based approaches.
Appendix N Additional Discussion
It is challenging to directly sort a sequence of generic data instances without using auxiliary networks and explicit supervision. Unlike earlier sorting methods, this sorting network-based research (Petersen et al., 2021; 2022) including our work ensures that we can train a neural network that predicts numerical scores and eventually sorts them, even though we do not necessitate accessing explicit supervision such as exact numerical values of the contents in high-dimensional data. In this sense, the practical significance of our proposed methods can be highlighted by offering this possibility of solving a sorting problem with high-dimensional inputs. For example, as shown in Section 5, we can compare images of street view house numbers using the sorting network where our neural network is trained without exact house numbers.
Moreover, instead of using costly supervision, our networks allow us to sort high-dimensional instances in a sequence where information on comparisons between instances is only given. This scenario often occurs when we cannot obtain complete supervision. For example, if we would sort four-digit MNIST images, ordinary neural networks are designed to solve a classification task by predicting class probabilities each of which indicates one of all labels from “$0$” to “$9$”. If some labels are missing and further we do not know the exact number of labels, they might fail in predicting unseen data corresponding to those labels. Unlike these methods, it is possible to solve sorting problems using our networks in such a scenario.
Furthermore, this study can be applied in diverse deep learning tasks for learning to sort generic high-dimensional data, such as information retrieval (Cao et al., 2007; Liu, 2009) and top-$k$ classification (Berrada et al., 2018).